EOQ, reorder point, and safety stock Alexis Company uses 631 units of a product per year on a continuous basis. The product has a fixed cost of $41 per order, and its carrying cost is $4 per unit per year. It takes 5 days to receive a shipment after an order is placed, and the firm wishes to hold 10 days usage in inventory as a safety stock. a. Calculate the EOQ. b. Determine the average level of inventory. (Note: Use a 365-day year to calculate daily usage.) c. Determine the reorder point. d. Indicate which of the following variables change if the firm does not hold the safety stock: (1) order cost, (2) carrying cost, (3) total inventory cost, (4) reorder point, (5) economic order quantity. a. Alexis EOQ is units. (Round to the nearest whole number.)
Answer #1

Answer a:

EOQ = SQRT (2 * Annual Demand in units * Order cost / Carrying Cost)

= SQRT (2 * 631 * $41 / $4)

= 113.73

EOQ = 113.73 or 114 units

Answer b:

Average Inventory = Minimum Level + EOQ/2 = 10 * 631/365 + 113.73 / 2 = 74.15 or 74 units

Answer c:

Reorder level or Ordering level = Safety stock + Lead time consumption

Safety stock =  (Annual Demand/365) x 10 days = 631 /365 *10 = 17.288 units

Lead time consumption = Lead time * Daily usage = 5 * 631/365 = 8.644 units

Reorder level or Ordering level = 17.288 + 8.644 = 25.93 units

Reorder level = 25.93 or 26 units

Answer d:

Variables that will change if firm does not hold safety stock:

(2) Carrying cost

(3) Total Inventory cost

(4) Reorder point

Explanation:

Carrying cost = Average inventory * carrying cost per unit.

Average inventory includes safety stock.

Total inventory costs includes carrying cost

Reorder point = Safety stock + Lead time consumption

If safety stock = 0, then Reorder point = Lead time consumption